OpenStudy (anonymous):

solve for x: 5x squared +3=8

OpenStudy (anonymous):

\[\sqrt{5x}+3=8\]

OpenStudy (lgbasallote):

\[\sqrt{5x} = 8 - 3\] \[\sqrt{5x} = 5\] square both sides \[5x = 25\] solve for x

OpenStudy (anonymous):

thanks